OpenStudy (lgbasallote):

Achilles and a tortoise were racing on a 1 km track. Achilles runs on a constant 8 m/s speed and the tortoise's speed is 0.8 m/s. Halfway through the track, Achilles decided to take a nap since the tortoise was so far behind. However, when he woke up, the tortoise was already near the finish line. Achilles started to run again with constant speed of 8 m/s, but he lost to the tortoise by 3 seconds. How long did Achilles sleep? Note: Let the tortoise and Achilles' speed be constant

OpenStudy (anonymous):

T(torto) = 625 + 625 = 1250s to Cover 1000m T(achilles) = 62.5 + 62.5 = 125s to cover 1000m Since the Achilles reached 3 Seconds Later .it lost by 3 seconds means 125 + x = 1253 x = 1128s

OpenStudy (lgbasallote):

i don't think that's right.....

OpenStudy (lgbasallote):

what's that 625 on the first line by the way?

OpenStudy (anonymous):

\[S=\frac{ d }{ t }\]

OpenStudy (lgbasallote):

i doubt this is that simple

OpenStudy (anonymous):

8(1253-x)=1000 x=1128 seconds Looks correct to me.

OpenStudy (lgbasallote):

1253 - x?

OpenStudy (lgbasallote):

you do know i asked how long Achilles slept and not how long Achilles was running...right?

OpenStudy (anonymous):

x is the time for which Achilles sleeps...... So, 1253-x gives how long he ran.

OpenStudy (lgbasallote):

hmm then that's not right

OpenStudy (anonymous):

WHY???

OpenStudy (lgbasallote):

like i said...it's not simple

OpenStudy (lgbasallote):

the question is...what's the right solution...

OpenStudy (anonymous):

@lgbasallote which part Do u think...i made a Mistake...

OpenStudy (lgbasallote):

why are you asking me? i'm the asker here...you tell me....

OpenStudy (lgbasallote):

because you think you're right?

OpenStudy (anonymous):

Toto took 1250s To cover 1000m.... right.?

OpenStudy (lgbasallote):

no need to explain to me your solution..i get it

OpenStudy (lgbasallote):

..it's the book that says it's wrong..not me

myininaya (myininaya):

Did he shout? You guys play nice please. People can disagree with answers. I don't find it rude to do so. I'm not saying anyone's answer is right here. I'm just saying people have a right to disagree but please be respectful about it.

myininaya (myininaya):

I will be deleting anything I find nasty here. Thanks.

myininaya (myininaya):

Keep in mind. Discussion is an awesome tool in learning. :)

OpenStudy (shubhamsrg):

* :P

OpenStudy (shubhamsrg):

though no one asked,,but i guess yahoo is right..

OpenStudy (lgbasallote):

i suppose one needs to get the speeds first...

ganeshie8 (ganeshie8):

me too i dont see a mistake in the solution

OpenStudy (shubhamsrg):

whats the ans in the book ?

OpenStudy (shubhamsrg):

and speeds are already given right ?

OpenStudy (lgbasallote):

ahh i think i get what the question is asking

OpenStudy (anonymous):

I got 1128. using a much more complex method.

OpenStudy (lgbasallote):

hmm maybe I can get an idea from that complex method

OpenStudy (lgbasallote):

this is a headscratcher

OpenStudy (anonymous):

Alright. Using your head. visualize the problem. A 1000 meter track. Tortoise and Achilles begin running. Achilles pulls ahead and half way takes a nap. Thats after 62.5 seconds. He's already covered 500 m. At that same time Achilles sleeps, Tortoise has reached 50 m. Now using that knowledge, we know it will take the Tortoise 1187.5 s to finish. It'll take Achilles 62.5 s. So subtract Achilles time from Tortoise's. Which gives you. 1,187.5 s - 62.5 = 1,125 s. Add three seconds. = 1,128 s

OpenStudy (lgbasallote):

so your answer is 1128 too huh

OpenStudy (anonymous):

Yep. It makes complete sense. He napped for 1,128 seconds. in order to finish 3 seconds after tortoise.

OpenStudy (lgbasallote):

what if the 3 seconds doesn't mean that Achilles finished 3 seconds later?

OpenStudy (anonymous):

Oh it does. Because at 1,125 s. Achilles finished right with tortoise. But adding 3 seconds means he finished 3 seconds later.

OpenStudy (lgbasallote):

i meant 3 seconds can mean that Achilles' running time is 3 more than the tortoise's

OpenStudy (anonymous):

No. It doesn't mean that here.

myininaya (myininaya):

@lgbasallote I agree with Yahoo's answer also. Like this is the way I thought of it:  Since like everyone is saying A finished 3 seconds after T. So that means A finished in 1253 seconds.